In a closed economy, where the households save 0.1% of their income and the only tax is 15% tax on income , what will be the result of a rise in government spending of £200000?


1
Expert's answer
2022-02-01T06:55:26-0500

The multiplier is:

m = 1/(1 - 0.999×(1 - 0.15)) = 6.63.

The result of a rise in government spending of £200000 is and increase in total income by 6.63×200000 = £1326000.